OpenStudy (anonymous):

Find the integral of arccos x /sqrt (1-x^2)dx

OpenStudy (anonymous):

let u=arccos x then du = -1/sqrt{1-x^2}dx , then -du=1/sqrt{1-x^2}dx integrate -u du =- u^2/2

OpenStudy (campbell_st):

\[\int\limits \arccos(x) / \sqrt{1 - x^2} dx = \int\limits \arccos(x) \times 1/\sqrt{1 - x^2} dx\] this is integration by substitution let u = arccos(x) \[du/dx = 1/\sqrt{1-x^2}\] then \[du = 1/\sqrt{1 - x^2} dx\] the problem can be rewritten as \[\int\limits u du = 1/2 u^2 + c\] replace u with arccos(x) gives \[\int\limits \arccos(x)/\sqrt{1 - x^2} dx = 1/2 (\arccos (x))^2 + c\]

OpenStudy (campbell_st):

oops forgot the negative